5. An accountant walks part of the way and takes the bus the remaining part of the way to work. The total distance from his house to work is 23 miles. He walks at the rate of 4 miles per hour while the bus travels at a speed of 35 miles per hour. The total trip time from his house to work is 2 hours. Calculate the distance the account travels by foot and the distance the accountant travels by bus
This is what I came up with so far:
Dw+Db=23
Dw/4 + Db/35=2
DW= Distance walked
DB=Distance bused

Answer by scott8148(6628)

good so far...

w = 23 - b

substituting ___ ((23 - b) / 4) + (b / 35) = 2

multiplying by 140 (LCD) ___ 805 - 35b + 4b = 280

adding 31b-280 ___ 525 = 31b ___ 525/31 = b

substituting ___ w = 23 - 525/31